# New phenomena arising from ${\mathcal{L}}$-invariants of modular forms

### v1

@@ 6.2. Loci with fixed reductions
$X_{k}^{\pm}(\overline{r})=\left\{{\mathcal{L}}\in{\mathbb{P}}^{1}(\overline{{{\mathbb{Q}}}}_{p})~{}\Big{|}~{}\begin{subarray}{c}\displaystyle\text{~{}there exists a $\overline{{{\mathbb{Z}}}}_{p}$-linear $\operatorname{Gal}(\overline{{{\mathbb{Q}}}}_{p}/{\mathbb{Q}}_{p})$-stable~{}}\\ \displaystyle\text{~{}lattice $T\subseteq V_{k,{\mathcal{L}}}^{\pm}$ such that$T/\mathfrak{m}_{\overline{{{\mathbb{Z}}}}_{p}}T\simeq\overline{V}$}\end{subarray}\right\}.$
@@
If $\overline{r}=\overline{r}_{0}$ there is nothing to prove. Otherwise, we assume $\overline{r}=\chi_{1}\oplus\chi_{2}$ for distinct characters $\chi_{j}$, ordered as above. The rest of the proof follows from a famous observation of Ribet [61, Section 2].
Suppose that ${\mathcal{L}}\in X_{k}^{\pm}(\overline{r})$. Since $k>2$, the representation $V_{k,{\mathcal{L}}}^{\pm}$ is irreducible and defined over some finite extension $E\supseteq{\mathbb{Q}}_{p}$. By Ribet’s [61, Proposition 2.1] there exists $\mathcal{O}_{E}$-linear and $\operatorname{Gal}(\overline{{{\mathbb{Q}}}}_{p}/{\mathbb{Q}}_{p})$-stable lattice $\Lambda\subseteq V_{k,{\mathcal{L}}}^{\pm}$ for which
@@
To define $R_{\overline{r}}^{k,\pm}$, Kisin first constructed (see [48]), for each $k$, a natural quotient $R_{\overline{r}}\twoheadrightarrow R_{\overline{r}}^{k,\mathrm{st}}$ that parametrizes semi-stable deformations of $\overline{r}$, with Hodge–Tate weights $10<k-$ 1 and determinant $\omega^{k-1}$. The semi-stable condition is partially encoded in Kisin’s ring by means of the inertial type. For comparison, we are taking the trivial inertial type. Let $\mathrm{X}_{\overline{r}}^{k,\mathrm{st}}=\operatorname{Spf}(R_{\overline{r}}^{k,\mathrm{st}})^{\mathrm{rig}}$ be the rigid generic fiber. The rigid space $\mathrm{X}_{\overline{r}}^{k,\mathrm{st}}$ is always reduced and equidimensional. Conceptually, $\mathrm{X}_{\overline{r}}^{k,\mathrm{st}}$ breaks up into a union of the crystalline locus and two loci that contain semi-stable and non-crystalline representations with one of the two signs $\pm$. For most $\overline{r}$, these three loci do not intersect. For $k>2$, the only case where intersection is possible is when $\overline{r}\simeq\mathrm{ind}(\omega_{2}^{k-1})$. In that case, the issue is that, for either $\pm$, we have
@@
The ring $R_{\overline{r}}^{k,\pm}$ picks out the semi-stable and non-crystalline deformations with a fixed sign $\pm$. To define it, one relies on extending the trivial inertial type to a Galois type, in the terminology of Rozensztajn [64, Definition 5.1.1(3)]. Specifically, let $W_{{\mathbb{Q}}_{p}}$ be the Weil group and denote $\tau_{\pm}=\operatorname{unr}(\pm 1)\oplus\operatorname{unr}(\pm p)$ (a smooth $W_{{\mathbb{Q}}_{p}}$-representation). Then, $R_{\overline{r}}^{k,\pm}$ is defined as the maximal reduced quotient of $R_{\overline{r}}^{k,\mathrm{st}}$ supported on the union of components in $\mathrm{Spec}(R_{\overline{r}}^{k,\mathrm{st}})$ that contain at least one semi-stable lift $r$ of $\overline{r}$ such that the associated Weil–Deligne representation $\mathrm{WD}(r)$ has non-trivial monodromy and Frobenius acting by $\tau_{\pm}$. See [63, Section 2.3.3].
@@
by [64, Theorem 5.3.1] and Section 7.6 of op. cit.. Rozensztajn also proves in [64, Corollary 5.3.2] that this makes $X_{k}^{\pm}(\overline{r})$ a standard subset of ${\mathbb{P}}^{1}(\overline{{{\mathbb{Q}}}}_{p})$. This means it is built as follows. You begin with either ${\mathbb{P}}^{1}(\overline{{{\mathbb{Q}}}}_{p})$ or a disc $v_{p}({\mathcal{L}}-x)>r$ defined by an open inequality. From that, you remove a finite union of discs $v_{p}({\mathcal{L}}-y)\geq s$ defined by closed inequalities. The outcome is called a connected standard subset. A finite union of connected standard subsets is called a standard subset ([64, Definition 3.3.1-3.3.2]).
@@ 7.1. Thresholds
(11) $-\frac{(p+1)k}{2(p-1)}+\frac{p+3}{2(p-1)}\leq v_{p}({\mathcal{L}}_{f}).$
@@
The argument of loc. cit. is based on analysis in $p$-adic Hodge theory, and theorems ([7, 13]) proven via a similar $p$-adic Hodge theory approach, in the crystalline cases, also produce bounds that are a factor away from realities. (Tight bounds, in crystalline cases and for regular eigenforms are proven in [52, Theorem 1.12]. For a purely local approach to tight bounds, using $p$-adic local Langlands, see [4, Theorem M].)
Let us summarize the discussion in terms of Section 6.2. First, $\overline{r}_{k,\infty}^{\pm}=\operatorname{ind}(\omega_{2}^{k-1})$, and therefore $X_{k}^{\pm}(\operatorname{ind}(\omega_{2}^{k-1}))$ always contains a component $X_{\infty}$ around ${\mathcal{L}}=\infty$. The theorem in [8] is a result proving $X_{\infty}$ contains a disc of some explicit radius. This implies a bound like (11) for regular eigenforms, and it makes improving the result of op. cit. an interesting project.
@@ 7.2. Sign dependence and Galois representations
Let $\operatorname{unr}(a)$ denote the unramified character taking value $a$. The first point is that if $\overline{r}$ is irregular, then $\overline{r}\simeq\overline{r}\otimes\operatorname{unr}(-1)$. The second point is that, given any ¯${\mathcal{L}}\in\overline{{{\mathbb{Q}}}}_{p}$, we also have
@@
producing ${\mathcal{L}}$-invariants with sign independent slopes. (An example of such a component would be a rational disc containing neither 00 or not $\infty$.)
@@
(See [75, Theorem 2].) Therefore, setting $k=2$ and $a_{p}(f)=\pm p^{\frac{k}{2}-1}=\pm 1$, we see that if $d_{2}^{\pm}(\overline{\rho})$ is non-zero, then $\overline{\rho}$ has the claimed shape. We have proven that there is at most one sign $\pm$ for which $d_{2}^{\pm}(\overline{\rho})\neq 0$, and we have identified the action of $\operatorname{Gal}(\overline{{{\mathbb{Q}}}}_{p}/{\mathbb{Q}}_{p})$ in that case.
@@
    In this case, we see $a_{p}(f)=\pm 1\bmod p$, By Ribet’s level-raising theorem (see [62, Theorem 1, point 3.]), we conclude that there exists a $p$-new eigenform lifting $\overline{\rho}$, which witnesses $d_{2}^{\pm}(\overline{\rho})\neq 0$.∎
@@ 7.3. Galois, II
$\overline{\rho}$-label $N$ $k$ LMFDB label local at $5$ representation
@@
$\overline{\rho}_{3}$ $9$ 4 9.4.a.a $\operatorname{ind}(\omega_{2}^{3})$
@@ 8.1. Gouvêa’s distribution conjecture
We add, given the extra context provided by Section 7, that the Liu–Truong–Xiao–Zhao paper also proves a rather strict law. Under their assumptions, there is an a priori containment $\mathbf{x}_{T}(\overline{\rho})\subseteq[0,1]$. See [52, Theorem 1.12].
@@
Then, the sets $\mathbf{y}_{T}^{\pm}$ are equidistributed on $[-1,0]$ for Lebesgue measure as $T\rightarrow\infty$.
@@
First, in Conjecture 8.2 we predict an end result supported on $[-1,0]$, while Conjecture 8.1 predicts an end result supported on $[0,1]$. If we replaced ${\mathcal{L}}_{f}$ by ${\mathcal{L}}_{f}^{-1}$ in our conjecture, this difference would go away. This replacement may have conceptual advantages, in terms of Section 6.2, as ${\mathcal{L}}^{-1}\approx 0$ is where the semi-stable, non-crystalline, loci in local deformation spaces intersect crystalline loci. But, leaving ${\mathcal{L}}_{f}$ the way it is also has the practical advantage that it reminds us that $v_{p}({\mathcal{L}}_{f})\leq 0$ almost all the time.
@@ 8.3. Scatter plots for distributional conjecture
Figure 16. Normalized $3$-adic slopes of ${\mathcal{L}}$-invariants in level $\Gamma_{0}(3)$ for eigenforms of weight $6\leq k\leq 508$. (Total data: 10752 eigenforms.)
Figure 17. Normalized $5$-adic slopes of ${\mathcal{L}}$-invariants in level $\Gamma_{0}(5)$ for eigenforms of weight $10\leq k\leq 472$. (Total data: 18560 eigenforms.)
Figure 18. Normalized $7$-adic slopes of ${\mathcal{L}}$-invariants in level $\Gamma_{0}(7)$ for eigenforms of weight $4\leq k\leq 290$. (Total data: 10511 eigenforms.)
@@
(22) $a_{p}(k)$ =±$=\pm p^{\frac{k_{0}}{2}-1}\left(1-\frac{{\mathcal{L}}_{f}}{2}\cdot(k-k_{0})\right)+O((k-k_{0})^{2}).$
@@
We illustrate the above idea with a numerical example. Take $p=3$ and $k=52$ in tame level $N=1$. Table 3 lists the slopes of eigenforms of level $\Gamma_{0}$ which appear in a sequence of weights $3$-adically approaching 52. The slopes in weight $52$ appear at the bottom of the table. We visually cut the table at weight $52+2\cdot 3^{12}$, as the listed data after that point is constant and matches the data in weight $52$. [^3] Actually the data presented here is computed via the ghost series as the weights which appear in this table are too large to effectively compute true slopes. Thus, this data is contingent on the ghost conjecture which is not yet known in this case.
@@
In weight $k_{0}=52$ we are only interested in the slope $125=\frac{k_{0}}{2}-$ 1 of the eight newforms. These are the fifth through the twelfth slopes, when ordered as in the table. Let’s focus our attention on the fifth slope in each weight in the table, which is the first bolded slope. As we move from the bottom of the table upwards, this fifth slope is 25 in weights $52+2\cdot 3^{j}$ for $j\geq 12$ and for smaller $j$ the slope decreases by 1 in each line in the table. The same phenomenon occurs for the sixth through eighth slopes but now the cutoff points after which the slopes begin to decrease (by one) is $j=10$, $7$ and $4$, respectively. This behavior suggests the analytic expansions in (22) of $a_{p}(k)$ corresponding to these newforms in weight 52 behaves like a linear function. Indeed, if $a_{p}(k)$ were linear, by (22), we would have
$a_{p}(k)$ =±$=\pm p^{\frac{k_{0}}{2}-1}\left(1-\frac{{\mathcal{L}}_{f}}{2}\cdot(k-k_{0})\right)$
@@
Note that under (23), the valuations of ${\mathcal{L}}$-invariants determine the cutoff points where the slopes begin to decrease in Table 3. This exactly matches the numerical data we collected on ${\mathcal{L}}$-invariants! Specifically, in our example of weight 52 and level $\Gamma_{0}(3)$, the valuations of the ${\mathcal{L}}$-invariants are $-4,-4,-7,-7,-10,-10,-12,-12$ which are exactly the cutoffs (doubled) we observed in the table above.
@@
(24) $s_{i+d_{k_{0}}}=\begin{cases}\displaystyle v_{p}({\mathcal{L}}_{f_{i}})+\frac{k_{0}-2}{2}+v_{p}(k-k_{0})&\text{if~{}}v_{p}(k-k_{0})<-v_{p}({\mathcal{L}}_{f_{i}})\\ \displaystyle\frac{k_{0}-2}{2}&\text{if~{}}v_{p}(k-k_{0})\geq-v_{p}({\mathcal{L}}_{f_{i}}),\\ \end{cases}$
@@
This equation is the key for us as it relates the valuation of ${\mathcal{L}}_{f}$ to slopes, and slopes can be more readily controlled. However, it is possible that no such $k$ exists satisfying (25). Indeed, if say $v_{p}({\mathcal{L}}_{f})=0$ then (25) clearly cannot be satisfied. Fortunately, our numerical data suggests that $v_{p}({\mathcal{L}}_{f})$ is typically quite negative and, in weight $k_{0}$, is very rarely larger than $-\log(k_{0})/\log(p)$. In what follows, we will assume that there exists a weight $k$ satisfying (25) and hope that this excludes only a thin set of ${\mathcal{L}}$-invariants.
@@
Where does this heuristic come from? Recall that Gouvêa’s distribution conjecture (Conjecture 8.1) predicts that the normalized slopes $\frac{p+1}{k}\cdot s_{i}$ are uniformly distributed between 00 and $1$ as $k\to\infty$. If we remove the $O(\log k)$-term in the above heuristic, it would be asserting that the $i$-th normalized slope equals $\frac{i}{d_{k}}$. In particular, these normalized slopes would be perfectly, evenly, placed between 00 and $1$ for every $k$. Such a claim is certainly too strong and thus we need some error term for which we choose $O(\log k)$. Heuristic 8.5 is thus a strengthening of Gouvea’s distribution conjecture. Moreover, this heuristic follows from the author’s ghost conjecture when $\overline{\rho}$ is regular (see [11, Theorem 3.1(a)]) with this $O(\log k)$ error term.
@@
Figure 20. $5$-adic slopes of ${\mathcal{L}}$-invariants in level $\Gamma_{0}(5)$ for $10\leq k\leq 472$ (in blue) versus the heuristic predictions for $10\leq k\leq 630$. Total real data: 18560 eigenforms. Total heuristic data: 33069 “eigenforms”.

### v2

@@ 6.2. Loci with fixed reductions
[…]
@@
If $\overline{r}=\overline{r}_{0}$ there is nothing to prove. Otherwise, we assume $\overline{r}=\chi_{1}\oplus\chi_{2}$ for distinct characters $\chi_{j}$, ordered as above. The rest of the proof follows from a famous observation of Ribet [62, Section 2].
Suppose that ${\mathcal{L}}\in X_{k}^{\pm}(\overline{r})$. Since $k>2$, the representation $V_{k,{\mathcal{L}}}^{\pm}$ is irreducible and defined over some finite extension $E\supseteq{\mathbb{Q}}_{p}$. By Ribet’s [62, Proposition 2.1] there exists $\mathcal{O}_{E}$-linear and $\operatorname{Gal}(\overline{{{\mathbb{Q}}}}_{p}/{\mathbb{Q}}_{p})$-stable lattice $\Lambda\subseteq V_{k,{\mathcal{L}}}^{\pm}$ for which
@@
To define $R_{\overline{r}}^{k,\pm}$, Kisin first constructed (see [48]), for each $k$, a natural quotient $R_{\overline{r}}\twoheadrightarrow R_{\overline{r}}^{k,\mathrm{st}}$ that parametrizes semi-stable deformations of $\overline{r}$, with Hodge–Tate weights 0<k10<k-1 and determinant $\omega^{k-1}$. The semi-stable condition is partially encoded in Kisin’s ring by means of the inertial type. For comparison, we are taking the trivial inertial type. Let $\mathrm{X}_{\overline{r}}^{k,\mathrm{st}}=\operatorname{Spf}(R_{\overline{r}}^{k,\mathrm{st}})^{\mathrm{rig}}$ be the rigid generic fiber. The rigid space $\mathrm{X}_{\overline{r}}^{k,\mathrm{st}}$ is always reduced and equidimensional. Conceptually, $\mathrm{X}_{\overline{r}}^{k,\mathrm{st}}$ breaks up into a union of the crystalline locus and two loci that contain semi-stable and non-crystalline representations with one of the two signs $\pm$. For most $\overline{r}$, these three loci do not intersect. For $k>2$, the only case where intersection is possible is when $\overline{r}\simeq\mathrm{ind}(\omega_{2}^{k-1})$. In that case, the issue is that, for either $\pm$, we have
@@
The ring $R_{\overline{r}}^{k,\pm}$ picks out the semi-stable and non-crystalline deformations with a fixed sign $\pm$. To define it, one relies on extending the trivial inertial type to a Galois type, in the terminology of Rozensztajn [65, Definition 5.1.1(3)]. Specifically, let $W_{{\mathbb{Q}}_{p}}$ be the Weil group and denote $\tau_{\pm}=\operatorname{unr}(\pm 1)\oplus\operatorname{unr}(\pm p)$ (a smooth $W_{{\mathbb{Q}}_{p}}$-representation). Then, $R_{\overline{r}}^{k,\pm}$ is defined as the maximal reduced quotient of $R_{\overline{r}}^{k,\mathrm{st}}$ supported on the union of components in $\mathrm{Spec}(R_{\overline{r}}^{k,\mathrm{st}})$ that contain at least one semi-stable lift $r$ of $\overline{r}$ such that the associated Weil–Deligne representation $\mathrm{WD}(r)$ has non-trivial monodromy and Frobenius acting by $\tau_{\pm}$. See [64, Section 2.3.3].
@@
by [65, Theorem 5.3.1] and Section 7.6 of op. cit.. Rozensztajn also proves in [65, Corollary 5.3.2] that this makes $X_{k}^{\pm}(\overline{r})$ a standard subset of ${\mathbb{P}}^{1}(\overline{{{\mathbb{Q}}}}_{p})$. This means it is built as follows. You begin with either ${\mathbb{P}}^{1}(\overline{{{\mathbb{Q}}}}_{p})$ or a disc $v_{p}({\mathcal{L}}-x)>r$ defined by an open inequality. From that, you remove a finite union of discs $v_{p}({\mathcal{L}}-y)\geq s$ defined by closed inequalities. The outcome is called a connected standard subset. A finite union of connected standard subsets is called a standard subset ([65, Definition 3.3.1-3.3.2]).
@@ 7.1. Thresholds
(11) $\frac{(p+1)k}{2(p-1)}+\frac{p+3}{2(p-1)}\leq v_{p}({\mathcal{L}}_{f}).$
@@
The argument of loc. cit. is based on analysis in $p$-adic Hodge theory, and theorems ([7, 13]) proven via a similar $p$-adic Hodge theory approach, in the crystalline cases, also produce bounds that are a factor away from realities. (Tight bounds, in crystalline cases and for regular eigenforms are proven in [51, Theorem 1.12]. For a purely local approach to tight bounds, using $p$-adic local Langlands, see [4, Theorem M].)
Let us summarize the discussion in terms of Section 6.2. First, $\overline{r}_{k,\infty}^{\pm}=\operatorname{ind}(\omega_{2}^{k-1})$, and therefore $X_{k}^{\pm}(\operatorname{ind}(\omega_{2}^{k-1}))$ always contains a component $X_{\infty}$ around $\mathcal{L}$$\infty$. The theorem in [8] is a result proving $X_{\infty}$ contains a disc of some explicit radius. This implies a bound like (11) for regular eigenforms, and it makes improving the result of op. cit. an interesting project.
@@ 7.2. Sign dependence and Galois representations
Let $\operatorname{unr}(a)$ denote the unramified character taking value $a$. The first point is that if $\overline{r}$ is irregular, then $\overline{r}\simeq\overline{r}\otimes\operatorname{unr}(-1)$. The second point is that, given any ${\mathcal{L}}\in\overline{{{\mathbb{Q}}}}_{p}$, we also have
@@
producing $\mathcal{L}$}-invariants with sign independent slopes. (An example of such a component would be a rational disc containing neither 0 or not $\infty$.)
@@
(See [76, Theorem 2].) Therefore, setting $k=2$ and $a_{p}(f)=\pm p^{\frac{k}{2}-1}=\pm 1$, we see that if $d_{2}^{\pm}(\overline{\rho})$ is non-zero, then $\overline{\rho}$ has the claimed shape. We have proven that there is at most one sign $\pm$ for which $d_{2}^{\pm}(\overline{\rho})\neq 0$, and we have identified the action of $\operatorname{Gal}(\overline{{{\mathbb{Q}}}}_{p}/{\mathbb{Q}}_{p})$ in that case.
@@
    In this case, we see $a_{p}(f)=\pm 1\bmod p$, By Ribet’s level-raising theorem (see [63, Theorem 1, point 3.]), we conclude that there exists a $p$-new eigenform lifting $\overline{\rho}$, which witnesses $d_{2}^{\pm}(\overline{\rho})\neq 0$.∎
@@ 7.3. Galois, II
$\overline{\rho}$-label NN $k$ LMFDB label local at $5$ representation
@@
$\overline{\rho}_{3}$ 99 4 9.4.a.a $\operatorname{ind}(\omega_{2}^{3})$
@@ 8.1. Gouvêa’s distribution conjecture
We add, given the extra context provided by Section 7, that the Liu–Truong–Xiao–Zhao paper also proves a rather strict law. Under their assumptions, there is an a priori containment $\mathbf{x}_{T}(\overline{\rho})\subseteq[0,1]$. See [51, Theorem 1.12].
@@
Then, the sets $\mathbf{y}_{T}^{\pm}$ are equidistributed on [1,0][-1,0] for Lebesgue measure as $T\rightarrow\infty$.
@@
First, in Conjecture 8.2 we predict an end result supported on [1,0][-1,0], while Conjecture 8.1 predicts an end result supported on $[0,1]$. If we replaced ${\mathcal{L}}_{f}$ by ${\mathcal{L}}_{f}^{-1}$ in our conjecture, this difference would go away. This replacement may have conceptual advantages, in terms of Section 6.2, as ${\mathcal{L}}^{-1}\approx 0$ is where the semi-stable, non-crystalline, loci in local deformation spaces intersect crystalline loci. But, leaving ${\mathcal{L}}_{f}$ the way it is also has the practical advantage that it reminds us that $v_{p}({\mathcal{L}}_{f})\leq 0$ almost all the time.
@@ 8.3. Scatter plots for distributional conjecture
Figure 16. Normalized 33-adic slopes of $\mathcal{L}$}-invariants in level $\Gamma_{0}(3)$ for eigenforms of weight $6\leq k\leq 508$. (Total data: 10752 eigenforms.)
Figure 17. Normalized 55-adic slopes of $\mathcal{L}$}-invariants in level $\Gamma_{0}(5)$ for eigenforms of weight $10\leq k\leq 472$. (Total data: 18560 eigenforms.)
Figure 18. Normalized 77-adic slopes of $\mathcal{L}$}-invariants in level $\Gamma_{0}(7)$ for eigenforms of weight $4\leq k\leq 290$. (Total data: 10511 eigenforms.)
@@
(22) $a_{p}(k)$ $=\pm p^{\frac{k_{0}}{2}-1}\left(1-\frac{{\mathcal{L}}_{f}}{2}\cdot(k-k_{0})\right)+O((k-k_{0})^{2}).$
@@
We illustrate the above idea with a numerical example. Take $p=3$ and $k=52$ in tame level $N=1$. Table 3 lists the slopes of eigenforms of level $\Gamma_{0}$ which appear in a sequence of weights $3$-adically approaching 52. The slopes in weight $52$ appear at the bottom of the table. We visually cut the table at weight $52+2\cdot 3^{12}$, as the listed data after that point is constant and matches the data in weight 5252.333Actually the data presented here is computed via the ghost series as the weights which appear in this table are too large to effectively compute true slopes. Thus, this data is contingent on the ghost conjecture which is not yet known in this case.
@@
In weight $k_{0}=52$ we are only interested in the slope $25=\frac{k_{0}}{2}-1$ of the eight newforms. These are the fifth through the twelfth slopes, when ordered as in the table. Let’s focus our attention on the fifth slope in each weight in the table, which is the first bolded slope. As we move from the bottom of the table upwards, this fifth slope is 25 in weights $52+2\cdot 3^{j}$ for $j\geq 12$ and for smaller $j$ the slope decreases by 1 in each line in the table. The same phenomenon occurs for the sixth through eighth slopes but now the cutoff points after which the slopes begin to decrease (by one) is $j=10$, 77 and $4$, respectively. This behavior suggests the analytic expansions in (22) of $a_{p}(k)$ corresponding to these newforms in weight 52 behaves like a linear function. Indeed, if $a_{p}(k)$ were linear, by (22), we would have
$a_{p}(k)$ $=\pm p^{\frac{k_{0}}{2}-1}\left(1-\frac{{\mathcal{L}}_{f}}{2}\cdot(k-k_{0})\right)$
@@
Note that under (23), the valuations of $\mathcal{L}$}-invariants determine the cutoff points where the slopes begin to decrease in Table 3. This exactly matches the numerical data we collected on $\mathcal{L}$}-invariants! Specifically, in our example of weight 52 and level $\Gamma_{0}(3)$, the valuations of the $\mathcal{L}$}-invariants are 4,4,7,7,10,10,12,12-4,-4,-7,-7,-10,-10,-12,-12 which are exactly the cutoffs (doubled) we observed in the table above.
@@
[…]
@@
This equation is the key for us as it relates the valuation of ${\mathcal{L}}_{f}$ to slopes, and slopes can be more readily controlled. However, it is possible that no such $k$ exists satisfying (25). Indeed, if say $v_{p}({\mathcal{L}}_{f})=0$ then (25) clearly cannot be satisfied. Fortunately, our numerical data suggests that $v_{p}({\mathcal{L}}_{f})$ is typically quite negative and, in weight $k_{0}$, is very rarely larger than $\log(k_{0})/\log(p)$. In what follows, we will assume that there exists a weight $k$ satisfying (25) and hope that this excludes only a thin set of $\mathcal{L}$}-invariants.
@@
Where does this heuristic come from? Recall that Gouvêa’s distribution conjecture (Conjecture 8.1) predicts that the normalized slopes $\frac{p+1}{k}\cdot s_{i}$ are uniformly distributed between 0 and $1$ as $k\to\infty$. If we remove the $O(\log k)$-term in the above heuristic, it would be asserting that the $i$-th normalized slope equals $\frac{i}{d_{k}}$. In particular, these normalized slopes would be perfectly, evenly, placed between 0 and $1$ for every $k$. Such a claim is certainly too strong and thus we need some error term for which we choose $O(\log k)$. Heuristic 8.5 is thus a strengthening of Gouvea’s distribution conjecture. Moreover, this heuristic follows from the author’s ghost conjecture when $\overline{\rho}$ is regular (see [11, Theorem 3.1(a)]) with this $O(\log k)$ error term.
@@
Figure 20. 55-adic slopes of $\mathcal{L}$}-invariants in level $\Gamma_{0}(5)$ for $10\leq k\leq 472$ (in blue) versus the heuristic predictions for $10\leq k\leq 630$. Total real data: 18560 eigenforms. Total heuristic data: 33069 “eigenforms”.